Solve for w.
w + 4/5 = 4 1/2


Answer by stanbon(75887)

Solve for w.
w + 4/5 = 4 1/2
-----
w = (9/2) - (4/5)
-----
w = (45-8)/10
w = 37/10
w = 3.7
---------------
Cheers,
Stan H.
-------------